OpenStudy (anonymous):

The first four terms of a sequence are shown below: 7, 4, 1, -2 Which of the following functions best defines this sequence?

OpenStudy (anonymous):

i will give a gold medal

OpenStudy (freckles):

well I would check to see if it was arithmetic sequence first : do you have term-previous term=same number like I mean are all the following equal: 4-7=? 1-4=? -2-1=?

OpenStudy (anonymous):

hold up let me put the answeres in

OpenStudy (anonymous):

it wont let me copy and paste them in

OpenStudy (anonymous):

anyone?

OpenStudy (freckles):

well I asked you a question and you haven't answered yet so I don't know if you mean me or someone else because you don't understand what I'm asking you?

OpenStudy (freckles):

I'm asking you if all the differences are the same above 4-7=? 1-4=? -2-1=?

OpenStudy (freckles):

if we can establish all of these are the same then we have an arithmetic sequence and that means we have a common difference

OpenStudy (freckles):

\[a_n=a_1+d(n-1)\] d represents the common difference if there is one So I'm still asking you the following questions: 4-7=? 1-4=? -2-1=?

OpenStudy (anonymous):

i

OpenStudy (anonymous):

i have no idea what im doing lol

OpenStudy (freckles):

So you haven't done subtraction yet?

OpenStudy (freckles):

 do you want to try doing 1-4=? and -2-1=?